# Tactile Perception in Upper Limb Prostheses:

### v1

@@ Abstract
Our research investigates vibrotactile perception in four prosthetic hands with distinct kinematics and mechanical characteristics. We found that rigid and simple socket-based prosthetic devices can transmit tactile information and surprisingly enable users to identify the stimulated finger with high reliability. This ability decreases with more advanced prosthetic hands with additional articulations and softer mechanics. We conducted experiments to understand the underlying mechanisms. We assessed a prosthetic user’s ability to discriminate finger contacts based on vibrations transmitted through the four prosthetic hands. We also performed numerical and mechanical vibration tests on the prostheses and used a machine learning classifier to identify the contacted finger. Our results show that simpler and rigid prosthetic hands facilitate contact discrimination (for instance, a user of a purely cosmetic hand can distinguish a contact on the index finger from other fingers with 83% accuracy), but all tested hands, including soft advanced ones, performed above chance level. Despite advanced hands reducing vibration transmission, a machine learning algorithm still exceeded human performance in discriminating finger contacts. These findings suggest the potential for enhancing vibrotactile feedback in advanced prosthetic hands and lay the groundwork for future integration of such feedback in prosthetic devices.
@@ I INTRODUCTION
The propagation of cutaneous feedback in the human hand is naturally related to high-frequency vibrations which propagate in the skin away from the original point of contact [10]. During tool-extended sensing, Miller et al. [11] proved that a rod mechanically transduces impact location into vibratory patterns decoded by the somatosensory system. The mechanoreceptors in the human hand transduce the vibratory cues into neural response patterns that preserve the location-specifying information [11]. Thus, the authors provide evidence of how handheld tools function as sensory extensions of the human’s body. Likewise, as pointed out by Farina et al. [9], a prosthetic user has natural sensory feedback in addition to vision and, depending on the type of prosthetic device, can perceive proprioceptive information from residual muscles, hear and feel motor actuation, and perceive vibrations conveyed through the prosthetic socket. Thus, for prosthetic users, vibrations patterns are important tactile information usually felt on the stump and partake in users’ strategies to compensate for the lack of sensory information [7, 12].
Despite the possibility of providing insight into prosthetic users’ baseline perceptions and the implications for prosthetic design, limited research has been conducted to investigate and characterize the natural transmission of tactile signals through prosthetic devices. An evaluation of the threshold level for vibrotactile cues detection of socket-prosthetic limbs with respect to bone-anchored prostheses is performed in [13]. The authors demonstrate that bone-anchored and socket-prosthetic limbs can activate stump-level receptors upon vibratory stimulation of the prosthetic thumb. As expected, bone-anchored prostheses promote vibration perception more than socket-based prosthetic devices. Furthermore, according to the recent study by Amoruso et al. [14], the existence of intrinsic somatosensory feedback by artificial body parts has been demonstrated to be crucial in enabling precise motor commands.
@@
In our experience, meaningful texture information is transmitted through socket-based prosthetic devices [15, 16]. During empirical observations, users exhibited the capability to discern minor tactile interactions on the digits of their prosthetic limb and, notably, to reconstruct the specific finger that had been contacted with a degree of reliability significantly exceeding random chance. We speculate vibrations to play an important role in this phenomenon. Indeed, studying the impact response analysis of a generic prosthetic-shaped system, different acceleration patterns arise when a contact cue stimulates different fingers (please see Fig.1 for impact response analysis results). These accelerations reach the other side of the prosthetic socket and may constitute a crucial component of prosthetic user baseline perception. However, considering the complexity of bionic hands, it is evident that vibration transmission of tactile cues differs depending on the point of contact and the shape of the prosthesis. Moreover, the mechanical characteristic of a given prosthesis is also expected to affect the propagation of vibrations and, therefore, the intensity and quantity of cues. The factors that we expect to affect such a phenomenon are the presence of articulation and the stiffness of the prosthetic hand. Thus, if the type of prosthesis demotes acceleration propagation, a possible enhancement of the natural vibration transmission would be beneficial for tactile feedback perception and for enabling prosthetic devices to function as a sensory extension. However, to the best of the authors’ knowledge, no study ever investigated in detail how and to what extent these baseline perceptions occur in socket-based prosthetic devices nor provided a quantitative assessment of the roles of the influencing factors.
@@
This study examines the perception and transmission of vibratory cues through four prosthetic hands with different mechanical characteristics (Fig. 2). Specifically, we compare hands in a range from very simple, with lower degrees of freedom (DoF) and rigid mechanisms, to more sophisticated devices with an increasing number of articulations and soft mechanical designs. We first report on a study with a prosthetic user, assessing her capability to discriminate which finger is touched based only on vibration patterns passively transmitted by different hands through the same socket to a user’s stump. Then, we quantitatively characterize the phenomenon of mechanical vibration transmission of the four prosthetic hands in terms of impact forces and socket-level vibration generated by impacting them with a test machine on the five fingers. To do so, we design and place a set of vibration sensors in the socket. Our aim is to correlate those data with user perceptual performance and the different hand designs. Finally, we examine the information content of the acceleration signals recorded at the stump level to determine whether acceleration data can be used to improve tactile perception and partially compensate for the sensory loss in the softer, more articulated hands. In other terms, we ask ourselves whether the signals retrieved by the vibration sensors on the socket contain enough information to enable softer, more articulated hands to the same level of discrimination that a user exhibits with simpler, rigid hands. The paper is structured as follows. The prosthetic hands are presented in Section II. Sections III-V includes the experimental methods. Results are presented and discussed in Section VI and Section VII. Section VIII concludes the paper.
@@ II Materials: Prosthetic Hands
The four hands exhibit an increasing complexity in terms of articulations and DoF. The Cosmetic hand is a passive hand mainly made of plastic with a silicone glove. The VariPlus is a tridigit myo-electrically controlled hand, with a single actuated degree of freedom opposing a thumb to two fingers, with two passive fingers and a five-finger glove. The three active fingers (index, middle, thumb) are rigid, made with metallic components, and mechanically coupled (i.e. close simultaneously), and the remaining two follow passively. The I-Limb is a myo-electrically controlled hand with metallic components and a soft silicone glove with thicker tips. The five fingers are individually powered. Each finger, except for the thumb, has a small mechanical play on the hand sagittal plane (see Fig. 2), while they are rigid in the other directions. The thumb has a transverse mechanical play. The SoftHand Pro is mainly made of plastic components with a silicone glove. A system of elastic ligaments connects the different phalanxes and makes the hand adaptable to all objects, especially in the sagittal hand plane. At the same time, each finger has a small mechanical play in the coronal plane. The mechanical structure of the fingers of the four test hands represents four of the main types of finger joints [18]: the rigid joint (found in the index, thumb and middle fingers of VP), the continuous joint (in CH, and ring and little fingers of VP), the flexible joint (in IL) and the dislocatable joint (in SH).
III Tactile Feedback Perception Experiment
@@
III-A Participant
One prosthetic user takes part in the experiments and gives her written, informed consent to participate. The participant (43 years old, female) is affected by limb agenesis and is used to wearing a cosmetic prosthesis even if having experience with myoelectric hands. The subject has no cognitive impairment that could affect the ability to follow the instructions of the study. Approval of all ethical and experimental procedures and protocols was granted by the Local Ethics Committee of Area Vasta Nord-Ovest (CEAVNO), Tuscany, Italy, under Protocol No. 7803.
III-B Experimental Setup
@@
The same reference socket is used for the four hands selected to ensure that the sole differences arise from the different hands used. In the preliminary experiment, the subject’s own cosmetic hand and the integrated socket are tested. The subject’s cosmetic hand is a CH hand without a quick disconnect wrist. Tactile cues are delivered with a small hammer () on the prosthesis fingertips, perpendicular to the coronal hand plane (see Fig.2). The experimenter underwent training to ensure consistent contacts delivery.
III-C Experimental Methods
A preliminary experiment with the participant’s cosmetic prosthesis is carried out to choose the proper number of contacts to deliver. The daily use of the prosthesis is expected to promote the recognition of the different fingers contacts. The goal is to reach a significant number of impacts before the loss of motivation and concentration. Based on the results obtained with the subject’s own prosthesis, the number of impacts to be delivered to each finger is set to 20.
Each experiment (preliminary experiment included) consists of a familiarization phase, a test phase, and rest phase. In the familiarization phase, the subject experiences five contacts on each finger without sound and visual insulation. During the test phase, the subject is isolated. Twenty contacts for each finger are performed with the small hammer. The order of the finger to be contacted is randomized to prevent the subject from guessing the next one. After each impact, the participant is asked to answer the question: “Which finger was contacted?”. The actual finger contacted and the corresponding answer from a user are saved for analysis. A rest phase of about fifteen minutes is done between one experiment and the other. The experiment is repeated four times, one for each of the prosthetic hands. The order of the specimen is randomized. The participant is informed that the experiment can be interrupted at any time. At the end of the experiment, the participant undergoes through a subjective evaluation procedure based on a brief questionnaire about prosthetic finger contact perception depending on the prosthetic hand used (Table I).
@@ III-C Experimental Methods
III-D Data Analysis
Participant’s prediction performance for each hand determines the four scores: (True Positives counts), (True Negatives counts), (False Positive counts) and (False Negative counts). Those scores are further analyzed using a confusion matrix, and the three metrics accuracy, recall, and precision [19]. Accuracy, defined as:
@@ III-D Data Analysis
where is the number of classes (thumb, index, middle, little, and ring). The accuracy [19] considering only the ability to discriminate the thumb finger with respect to the others and the index finger compared to the other fingers are also computed.
IV Tactile Feedback Transmission Experiment
@@ IV-A Experimental Setup
The four hands are mounted on the reference socket. The same supporting structure as in Section III is used. We attach five IMUs (MPU-9250) to the inner surface of the socket in a radial distribution to measure the differences in vibration patterns after each contact (Fig. 4, 5). This choice is grounded upon the outcomes of the impact analysis performed on a generic prosthetic-shaped system modelled as a continuous body (as shown in Fig. 1). Although the analysis simplifies the system compared to a real prosthetic system, it demonstrates that the same contact on each finger generates distinct acceleration patterns that reach the other side of the prosthesis. Furthermore, IMUs’ position and distribution are also justified by the physical constraints of the residual limb of the participant and its interfacing with the socket.
@@
A sensorized pendulum (see Fig. 4) is used by the experimenter to contact the prostheses fingertips with repeatable impacts. As in the first experiment, each contact is perpendicular to the coronal hand plane (Fig. 2). The pendulum can be moved and adjusted to contact a particular finger in a given position. Before each impact, the pendulum arm is aligned to an adjustable end-stop to hold it at a desired angle. Then, the pendulum is released to impact the prosthesis. During each impact, a load cell (ATI Nano17 Transducer) at the lower extremity of the pendulum measures the impact forces while an encoder (AS5045) at the upper extremity the angles.
IV-B Experimental Method
@@
In preliminary tests, the pendulum angle that can create an impact without reaching IMUs full scale is determined and kept for all experiments (). During impacts, acceleration signals inside the socket and impact forces at each fingertip are recorded through electronic boards [20]. A custom C++ software is developed to synchronize, start, and stop the recordings. The experiment starts when the pendulum is set at the predetermined angle through the encoder and the end-stop. The pendulum is left free to swing, and the load cell surface contacts the artificial fingertip. After each contact, the prosthetic limb oscillates. From the load cell, it is possible to record the three components of the impact force , while from each IMU inside the socket, the three components of the acceleration . The recording session is done at to ensure that each impact is properly captured and within the limits of the hardware used to register. The experiment ends when the pendulum is manually stopped after adequately completing the impact. Since the recorded signals have proven to be repeatable, three strikes for each finger are performed, and the best one is saved for the subsequent analysis. The experiment is repeated for each robotic hand.
IV-C Data Analysis
@@
After the experimental session, all signals are analyzed and processed. The impact force component is selected as the main contribution of the contacts because the pendulum swings in the same direction. As regards acceleration signals, various processes are carried out to analyze high-frequency signals only. Outliers are removed from the three components . Then, a high-pass filter is used to remove the low-frequency oscillation of the prosthesis caused by the contacts. Forces and high-frequency accelerations are cut to isolate the contact. Each acceleration signal is cut with respect to the impact force peak. The contact is accurately represented by three hundred samples (). The next step is assigning an energy value to each signal to compare each bionic hand’s results quantitatively. With the algorithm DFT321 [21], the three components are then reduced into a one-dimensional signal with the same energy as their sum (Fig. 6). The DTF321, as part of the algorithms that seek the spectral difference, is considered by Lee at al. [22] one of the best 321 approaches for offline processing when perceptual similarity is prioritized.
@@
V Tactile Feedback Recognition Experiment

### v2

@@ Abstract
Tactile feedback is essential for upper-limb prostheses functionality and embodiment, yet its practical implementation presents challenges. Users must adapt to non-physiological signals, increasing cognitive load. However, some prosthetic devices transmit tactile information through socket vibrations, even to untrained individuals. Our experiments validated this observation, demonstrating a user’s surprising ability to identify contacted fingers with a purely passive, cosmetic hand. Further experiments with advanced soft articulated hands revealed decreased performance in tactile information relayed by socket vibrations as hand complexity increased. To understand the underlying mechanisms, we conducted numerical and mechanical vibration tests on four prostheses of varying complexity. Additionally, a machine-learning classifier identified the contacted finger based on measured socket signals. Quantitative results confirmed that rigid hands facilitated contact discrimination, achieving 83% accuracy in distinguishing index finger contacts from others. While human discrimination decreased with advanced hands, machine learning surpassed human performance. These findings suggest that rigid prostheses provide natural vibration transmission, potentially reducing the need for tactile feedback devices, which advanced hands may require. Nonetheless, the possibility of machine learning algorithms outperforming human discrimination indicates potential to enhance socket vibrations through active sensing and actuation, bridging the gap in vibration-transmitted tactile discrimination between rigid and advanced hands.
@@ I INTRODUCTION
Limited research has explored the natural transmission of tactile signals through prosthetic devices despite its potential to offer insights into users’ baseline perceptions and ultimately contribute to the refinement and optimization of haptic feedback systems. During tool-extended sensing, Miller et al. [10] proved that a rod mechanically transduces impact location into vibratory patterns decoded by the somatosensory system. The mechanoreceptors in the human hand transduce the vibratory cues into neural response patterns that preserve the location-specifying information [10]. Thus, the authors provide evidence of how handheld tools function as sensory extensions of the human body. Likewise, as pointed out by Farina et al. [9], a prosthetic user has natural sensory feedback in addition to vision and, depending on the type of prosthetic device, can hear and feel motor actuation, and perceive vibrations conveyed through the prosthetic socket. An evaluation of the threshold level for contact cues detection of socket-prosthetic limbs with respect to bone-anchored prostheses is performed in [11]. The authors demonstrate that bone-anchored and socket-prosthetic limbs can activate stump-level receptors upon vibratory stimulation of the prosthetic thumb. Furthermore, according to the recent study by Amoruso et al. [12], the existence of intrinsic somatosensory feedback by artificial body parts has been demonstrated to be crucial in enabling precise motor commands. Thus, for prosthetic users, vibration patterns are important tactile information usually felt on the stump and partake in users’ strategies to compensate for the lack of sensory information [7, 13].
@@
We posit that investigating natural tactile information transmission is crucial for developing intuitive, non-invasive haptic prosthetic devices, aiming to optimize conveyed information and reduce users’ cognitive load. In our experience, meaningful texture information is transmitted through socket-based prosthetic devices [14, 15]. During empirical observations, users exhibited the capability to discern minor tactile interactions on the digits of their prosthetic limb and, notably, to reconstruct the specific finger that had been contacted with a degree of reliability significantly exceeding random chance. Considering the complexity of bionic hands, it is evident that the transmission of tactile cues differs depending on the point of contact and the shape of the prosthesis. Moreover, the mechanical characteristic of a given prosthesis is also expected to affect vibration propagation and, therefore, the intensity and quantity of cues. The factors we expect to influence this phenomenon include the prosthesis design, the presence of articulations, the rigidity of the assembled mechanisms, and the stiffness of the materials used. To the best of the authors’ knowledge, no study ever investigated how and to what extent these baseline perceptions occur in socket-based prosthetic devices nor provided a quantitative assessment of the roles of the influencing factors.
This study examines the perception and transmission of tactile cues through four prosthetic hands with different characteristics (Fig. 1).
@@
Specifically, we compare hands ranging from rigid, characterized by lower degrees of freedom (DoF), rigid mechanisms, and stiff materials, to advanced ones with more articulations, flexible mechanical designs, and softer materials. We first report on a pilot study with a prosthetic user, assessing their capability to discriminate which finger is contacted based only on vibration patterns passively transmitted by different hands through the same socket to a user’s stump. Then, we quantitatively characterize the phenomenon of mechanical vibration transmission of the four prosthetic hands in terms of impact forces and socket-level vibration generated by impacting them with a test machine on the five fingers. To do so, we design and place a set of vibration sensors in the socket. Our aim is to correlate those data with user perceptual performance and the different hand designs. Finally, we examine the information content of the acceleration signals recorded at the stump level to determine whether acceleration data can be used to improve tactile perception and partially compensate for the sensory loss in the softer, more articulated hands. In other terms, we ask ourselves whether the signals retrieved by the vibration sensors on the socket contain enough information to enable more advanced hands to the same level of discrimination that a user exhibits with rigid hands.
@@ II Materials: Prosthetic Hands
The four hands range from rigid designs with fewer DoF, rigid mechanisms, and stiff materials to advanced designs with more articulations, flexible designs, and softer materials. The Cosmetic hand is a passive hand mainly made of plastic with a silicone glove. The VariPlus is a tridigit myo-electrically controlled hand with a single actuated degree of freedom opposing a thumb to two fingers, with two passive fingers and a five-finger glove. The three active fingers (index, middle, thumb) are rigid, made with metallic components, and mechanically coupled (i.e. close simultaneously), and the remaining two follow passively. The I-Limb is a myo-electrically controlled hand with metallic components and a soft silicone glove with thicker tips. The five fingers are individually powered. Each finger, except for the thumb, has a small mechanical play on the hand sagittal plane (see Fig. 1), while they are rigid in the other directions. The thumb has a transverse mechanical play. The SoftHand Pro is mainly made of plastic components with a silicone glove. A system of elastic ligaments connects the different phalanxes and makes the hand adaptable to all objects, especially in the sagittal hand plane. At the same time, each finger has a small mechanical play in the coronal plane. The mechanical structure of the fingers of the four test hands represents four of the main types of finger joints [17]: the rigid joint (found in the index, thumb and middle fingers of VP), the continuous joint (in CH, and ring and little fingers of VP), the flexible joint (in IL) and the dislocatable joint (in SH).
III Perliminary Analysis
We speculate vibrations to play an important role in the natural tactile information transmission through socket-based prosthetic devices. As a preliminary exploration of this phenomenon, we conduct a response analysis of a simplified generic prosthetic-shaped system. The system is modelled as a continuous body consisting of a socket and hand plastic model (), resembling a typical cosmetic prosthesis. Our investigation revealed that different acceleration patterns arise when a contact cue stimulates different fingers, with each finger experiencing a force of 2N at its respective fingertip (please refer to Fig.2).
These accelerations reach the other side of the prosthetic socket and may constitute a crucial component of prosthetic user baseline perception. These findings have paved the way for subsequent experiments, motivating our study and prompting further investigation.
IV Tactile Feedback Perception Experiment
@@
IV-A Participant
One prosthetic user takes part in the experiments and gives their written, informed consent to participate. The participant (43 years old, female) is affected by limb agenesis and is used to wearing a cosmetic prosthesis even if having experience with myoelectric hands (a SoftHand Pro and a MyoHand VariPlus Speed device). The participant’s limited experience with the latter devices suggests minimal potential impact on the results. The subject has no cognitive impairment that could affect the ability to follow the instructions of the study. Approval of all ethical and experimental procedures and protocols was granted by the Local Ethics Committee of Area Vasta Nord-Ovest (CEAVNO), Tuscany, Italy, under Protocol No. 7803.
IV-B Experimental Setup
@@
The same reference socket is employed for all four selected prosthetic hands to ensure that differences arise solely from the variations in hands used. This Ottobock444Ottobock, [Online], Available: https://www.ottobock.com/en-us/home socket, crafted from carbon prepreg for the outer component and silicone for the inner component, was specifically customized for the participant’s use. This choice was made to facilitate seamless compatibility with all four prosthetic hands utilized in the study and to replicate the standard socket configuration typically associated with myoelectric prostheses. The participant’s personal prosthesis consists of a cosmetic hand identical to the Ottobock1 CH hand used in this study, without a quick disconnect wrist, and integrated with a plastic and silicone cosmetic socket to create a unified structure. This personal prosthesis does not include any EMG sensors. In the preliminary experiment, the subject’s personal cosmetic prosthesis is tested. Tactile cues are delivered with a small hammer () on the prosthesis fingertips, perpendicular to the coronal hand plane (see Fig.1). The experimenter underwent training to ensure consistent contacts delivery. The training involved repeated practice sessions to ensure consistent impact delivery to each finger in terms of intensity and direction. Despite diligent efforts, variations in execution were recognized during experimental sessions.
IV-C Experimental Methods
A preliminary experiment with the participant’s cosmetic prosthesis is carried out to determine the optimal number of impacts for assessing the subject’s recognition performance while also ensuring they did not experience undue fatigue or a loss of motivation or concentration. In order to optimize the subject’s comfort, we examined different approaches for delivering impacts, including positioning the subject with their arm suspended or the end of their arm resting on the edge of a table. Ultimately, we select the most suitable position for the user, as depicted in Fig. 3, with the prosthesis hanging from a supporting structure. We then gradually increase the number of impacts per finger, starting with a threshold of 5 impacts. The subject tolerated up to 20 impacts per finger without experiencing discomfort or fatigue. Based on these findings, we established that delivering 20 impacts to each finger would be appropriate for our experimental purposes.
Each experiment (preliminary experiment included) consists of a familiarization phase, a test phase, and rest phase. In the familiarization phase, the subject experiences five contacts on each finger without sound and visual insulation. During the test phase, the subject is isolated. Twenty contacts for each finger are performed with the small hammer. The order of the finger to be contacted is randomized to prevent the subject from guessing the next one. After each impact, the participant is asked to answer the question: “Which finger was contacted?”. The actual finger contacted and the corresponding answer from a user are saved for analysis. A rest phase of about fifteen minutes is done between one experiment and the other. The experiment is repeated four times, one for each of the prosthetic hands. The order of the specimens is randomized (IL, VP, CH, SH). The participant is informed that the experiment can be interrupted at any time. At the end of the experiment, the participant undergoes a subjective evaluation procedure based on a brief questionnaire about prosthetic finger contact perception depending on the prosthetic hand used (Table I).
@@ IV-C Experimental Methods
IV-D Data Analysis
Participant’s prediction performance for each hand determines the four scores: (True Positive counts), (True Negative counts), (False Positive counts) and (False Negative counts). Those scores are further analyzed using a confusion matrix, and the three metrics accuracy, recall, and precision [18]. Accuracy, defined as:
@@ IV-D Data Analysis
where is the number of classes (thumb, index, middle, little, and ring). The accuracy [18] considering only the ability to discriminate the thumb finger with respect to the others and the index finger compared to the other fingers are also computed. This assessment holds significance due to the pivotal roles of these fingers in tasks requiring precise grasping motions. It underscores the necessity of evaluating their discriminative ability, as it directly impacts the efficiency and precision of prosthetic users in performing daily activities.
V Tactile Feedback Transmission Experiment
@@ V-A Experimental Setup
The four hands are mounted on the reference socket. The same supporting structure as in Section IV is used. We attach five IMUs (MPU-9250) to the inner surface of the socket in a radial distribution to measure the differences in vibration patterns after each contact (Fig. 4, 5). This choice is grounded upon the outcomes of the impact analysis performed on a generic prosthetic-shaped system modelled as a continuous body (as shown in Fig. 2). Although the analysis simplifies the system compared to a real prosthetic system, it demonstrates that the same contact on each finger generates distinct acceleration patterns that reach the other side of the prosthesis. Furthermore, IMUs’ position and distribution are also justified by the physical constraints of the residual limb of the participant and its interfacing with the socket.
@@
A sensorized pendulum (see Fig. 4) is used by the experimenter to contact the prosthesis fingertips with repeatable impacts. As in the first experiment, each contact is perpendicular to the coronal hand plane (Fig. 1). The pendulum can be moved and adjusted to contact a particular finger in a given position. Before each impact, the pendulum arm is aligned to an adjustable end-stop to hold it at a desired angle. Then, the pendulum is released to impact the prosthesis. During each impact, a load cell (ATI Nano17 Transducer) at the lower extremity of the pendulum measures the impact forces while an encoder (AS5045) at the upper extremity the angles.
V-B Experimental Method
@@
In preliminary tests, the pendulum angle that can create an impact without reaching IMUs full scale is determined and kept for all experiments (). During impacts, acceleration signals inside the socket and impact forces at each fingertip are recorded through electronic boards [19]. A custom C++ software is developed to synchronize, start, and stop the recordings. The experiment starts when the pendulum is set at the predetermined angle through the encoder and the end-stop. The pendulum is left free to swing, and the load cell surface contacts the artificial fingertip. After each contact, the prosthetic limb oscillates. From the load cell, it is possible to record the three components of the impact force , while from each IMU inside the socket, the three components of the acceleration . The recording session is done at to ensure that each impact is properly captured and within the limits of the hardware used to register. The experiment ends when the pendulum is manually stopped after adequately completing the impact. Since the recorded signals have proven to be repeatable, three strikes per finger are performed, and one is saved for analysis. This avoids averaging, ensures realistic signals, leverages repeatability, and prioritizes signals without data loss; if no loss occurs, one is chosen randomly. The experiment is repeated for each robotic hand.
V-C Data Analysis
@@
After the experimental session, all signals are analyzed and processed. The impact force component is selected as the main contribution of the contacts because the pendulum swings in the same direction. In regards to the acceleration signals, various processes are carried out to analyze high-frequency signals only. Communication errors between the electronic boards and the IMUs are removed from the three components . Then, a high-pass filter is used to remove the low-frequency oscillation of the prosthesis caused by the contacts. Forces and high-frequency accelerations are cut to isolate the contact. Each acceleration signal is cut with respect to the impact force peak. The contact is accurately represented by three hundred samples (). The next step is assigning an energy value to each signal to compare each bionic hand’s results quantitatively. With the algorithm DFT321 [20], the three components are then reduced into a one-dimensional signal with the same energy as their sum (Fig. 6). The DTF321 algorithm, which aims to identify spectral differences, was utilized in our previous works [15, 21].
@@
VI Tactile Feedback Recognition Experiment